Which of the following are measurements of the sides of a right triangle?
a) 25,23,7
b) 9,6,3
c) 18,15,4
d) none of the above

Answer by stanbon(75887)

Which of the following are measurements of the sides of a right triangle?
Check c^2 = a^2 + b^2 where c is the largest of the 3 sides.
If c^2 = a^2+b^2 you have a right triangle. If it doesn't, you don't.
-------------------
Cheers,
Stan H.
==================
a) 25,23,7
b) 9,6,3
c) 18,15,4